OpenStudy (anonymous):

A manufacturer of mountain bikes has found that when 20 bikes are produced per day, the average cost is $200 and the marginal cost is $150. Based on the information, approximate the total cost of producing 21 bikes per day.

OpenStudy (slaaibak):

If the average cost is 200 per 20 bikes, the total cost for 20 bikes is: 200 * 20 Then you have to add another bike, which has a marginal cost of 150 so it's 200 * 20 + 150

OpenStudy (anonymous):

The manufacturer of a product has found that when 20 units are produced per day, the average cost is $150 and the marginal cost is $125. What is the relative rate of change of the average cost with respect to quantity when q=20? How would i set up the equation? I know how to differentiate and find relative rate of change... but I get lost when trying to generate an equation.

OpenStudy (slaaibak):

Hmmm. Total cost = Average Cost * Quantity Marginal cost (Change in total cost) = Average cost * Change in quanity + Change in average * Quanity 125 = 200 * 0 + 20 * Change in average cost 125/20 = change in average cost. Actually, I have no Idea how to do this, I did economics a year ago. But maybe this helps?

OpenStudy (slaaibak):

Although this should be negative, since the average cost will decrease when the marginal cost is lower than the average cost.

OpenStudy (anonymous):

the answer in the book says -1/120

OpenStudy (anonymous):

I agree, it's just frustrating trying to generate an equation :P but the relationships helped, thanks :)

OpenStudy (slaaibak):

I disagree with your book. -1/120 is way too small :/

OpenStudy (anonymous):

maybe it is wrong, i am more concerned with figuring out how to write the equation... math exam in a few weeks!

OpenStudy (anonymous):

average cost per unit: C = c/q where c is cost and q is quantity

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

if you isolate for c (cost), you get Cq = c or 150*20 = 3000?

OpenStudy (slaaibak):

Everything starts with this formula: TC = AC * Q Deriving this: dTC/dQ = dAC/dQ * Q + A(dQ/dQ) dTc/dQ = Marginal cost Marginal cost = Q * dAc/dQ + A 125 = 20 * dAc/dQ + 150 -25 = 20 * dAc/dQ -25/20 = dAC/dQ So the answer is -5/4

OpenStudy (anonymous):

total cost = average cost * quantity? just to clarify

OpenStudy (slaaibak):

Yes

OpenStudy (anonymous):

and what is A?

OpenStudy (slaaibak):

Sorry, meant Average cost

OpenStudy (anonymous):

perfect, i will try to work through this using that formula, and consult my TA about it when we return to classes on Monday :) thanks so much

OpenStudy (slaaibak):

Cool, no problem. Just remember, try writing the basic formulas and then deriving those. Key points to know: Total cost = Average Cost * Quantity Marginal cost = Change in total cost = dTC/dQ